OpenStudy (shelby1290):

what amount of energy does a kettle use to boil water if it has 800C of charge passing through it with a potential difference of 120V?

OpenStudy (shelby1290):

would the formula change to charge=difference in potential difference / potential difference so 800/120 which is 6.6=7 amps?

OpenStudy (wolfe8):

I don't understand. I thought you are looking for the energy

OpenStudy (shelby1290):

oh sorry i thought i had to switch the formula for potential difference around

OpenStudy (shelby1290):

so 800 divided by 120?

OpenStudy (shelby1290):

following E= Q/V

OpenStudy (wolfe8):

Oh no they are multiplied together. E=QV=Q*V

OpenStudy (shelby1290):

SORRY

OpenStudy (shelby1290):

so it would be 800x120

OpenStudy (shelby1290):

which would be 96000

OpenStudy (wolfe8):

Yes

OpenStudy (shelby1290):

joules?

OpenStudy (shelby1290):

or C

OpenStudy (shelby1290):

oops A

OpenStudy (wolfe8):

Yes Joules since it is C times J/C. (V=J/C)

OpenStudy (shelby1290):

oh okay thank you

OpenStudy (wolfe8):

You're welcome :)